LASER EXCITATION AND DISPERSED FLUORESCENCE SPECTRA OF PtC_{2}

\underline{ROBERT R. BOUSQUET}, HIYUN KIM, TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287.

The laser excitation spectrum of molecules produced in a molecular beam formed by laser ablation of a platinum rod in the presence of a methane/argon mixture in a supersonic free-jet expansion has been recorded. A series of bands observed at 18200, 18440 and 18660 cm^{−1} is attributed to PtC_{2}. The molecule is proposed to have a T-shaped structure with C_{2}v symmetry, analogous to YC_{2} ^{a}. The dispersed fluorescence spectra observed by pumping the bands show progressions in both ν_{2} (a_{1}, Pt-(C_{2}) st.) and ν_{3} (b_{2}, Pt-(C_{2}) rock) in addition to several bands involving combinations of these modes. Preliminary analysis has yielded frequencies ν′′_{2} = 485 cm^{−1} and ν′′_{3} = 223 cm^{−1}. Isotopic substitution spectra have also been recorded using carbon-13 substituted methane in the carbon source. A discussion of the vibrational analysis will be given.

aT. C. Steimle, A. J. Marr, J. Xin, A. J. Merer, K. Athanassenas and D. Gillett, J. Chem. Phys. 106, 2060-2066 (1997)